# ECM: A Unified Electronic Circuit Model for Explaining the Emergence of In-Context Learning and Chain-of-Thought in Large Language Model

### v1

@@ B.1.2 Analysis of Effectiveness of Coverage Design
Similarly, we hypothesize that coverage can be considered correct if at least one of multiple parallel reasoning is accurate. Under this assumption, the integration difficulty in self-consistency verification approaches zero. Analogously, this can be modeled as a zero-value combined resistance in parallel with multiple CoT resistors, where the coverage resistance of self-consistency $\hat{R}_{\text{all}}$ is given by:
$\hat{R}_{\text{all}}=R_{0}+R_{S}+\frac{1}{\sum_{i=1}^{n}\frac{1}{R_{\text{CoT}}^{i}}}=R_{0}+\frac{1}{\sum_{i=1}^{n}\frac{1}{R_{\text{CoT}}^{i}}}$ (17)
@@
$P^{\text{cover}}_{\text{out}}=\frac{\left(\mathcal{E}_{\text{model}}+\mathcal{E}_{\text{ICL}}\right)^{2}R_{0}}{(\frac{1}{\sum_{n}\frac{1}{R_{\text{CoT}}}}+R_{0})^{2}},$ (18)
which are much more smaller than $P^{\text{self}}_{\text{out}}$ in the same $n$ sample size with $R_{S}\gg 0$. As the number of samples increases, the curve in Figure 4c demonstrates that the real performance aligns closely with the theoretical range. Building upon the same bounded assumption of $\{R_{\text{CoT}}^{i}\}_{i=1}^{\infty}$, this setup also defines a resistance lower bound:
$\lim_{n\rightarrow+\infty}\hat{R}_{\text{all}}=R_{0}+\lim_{n\rightarrow+\infty}\frac{1}{\sum_{i=1}^{n}\frac{1}{R_{\text{CoT}}^{i}}}=R_{0}$ (19)
@@
$\hat{P}^{\text{cover}}_{\text{out}}=\frac{\left(\mathcal{E}_{\text{model}}+\mathcal{E}_{\text{ICL}}\right)^{2}}{R_{0}}.$ (20)
As shown in Figure 4c, the model’s performance gradually converges toward this upper bound and much larger than the performance of self-consistency, demonstrating consistency with the theoretical predictions.
@@
According to the theoretical formulation in Eq. 18, in the reasoning phase where the resistance is significantly greater than $R_{0}$, the overall trend shows that Coverage is approximately proportional to the number of samples. Meanwhile, certain Inference Scaling Laws indicate that Coverage and the logarithm of the number of samples exhibit an approximately proportional relationship. These two observations are not contradictory, as our model assumes relatively independent sampling. However, we have observed that when the number of samples exceeds 100, an increasing proportion of the samples become duplicates rather than independent samples. Consequently, the probability of generating new samples follows the trend where $n$ samples are expected to yield approximately $\log n$ distinct reasoning samples.
@@ B.2 Analysis of Effectiveness of Fine-grained Self-consistency
$R^{\prime}_{\text{all}}=R_{0}+R_{S}+\frac{1}{\sum_{i=1}^{n}\frac{1}{R_{\text{CoT}}^{i}}}.$ (21)
As shown in Figure 6a, this process is decomposed into multiple sequential, smaller-scale parallel reasoning and verifications. This refinement transforms a single large parallel reasoning resistance with significant verification resistance into a sequence of smaller parallel reasoning resistance and smaller verification resistance. Formally, this is expressed as:
@@
where $n$ denotes the sampling size, $R_{j}^{S}$ represents the small and fine-grained verification resistance for $j$-th step, and $R_{j}^{i}$ corresponds to smaller reasoning resistance for $j$-th step for $i$-th sampling. The total resistance can be broken down into resistance of sub-step resistor as $R^{i}_{\text{CoT}}=\sum_{j}R_{j}^{i}$.
@@
$R^{\text{fine}}_{\text{all}}-R^{\prime}_{\text{all}}$ $=\sum_{j}R_{j}^{S}-R_{S}+\frac{R_{\text{CoT}}}{n}-\sum_{j}\frac{R_{j}^{i}}{n}$ (23)
@@ B.3 Analysis of Effectiveness of Chain-of-Self-consistency
$R^{\text{CoV}}_{\text{all}}$ $=R_{0}+\frac{1}{\sum_{j=1}^{k}\sum_{i=1}^{n}\frac{1}{R_{\text{S}}}}+\sum_{j=1}^{k}\left(R_{\text{meta}}\right)+\frac{R_{\text{CoT}}}{n}$
$=R_{0}+\frac{R_{\text{S}}}{nk}+kR_{\text{meta}}+\frac{R_{\text{CoT}}}{n}.$
Clearly, as the sample size $n$ and the meta-validation steps $k$ increase, $\frac{R_{\text{S}}}{nk}$ decreases, leading to reduced difficulty in meta-validation. This significantly optimizes the total resistance and improves output power. Figure 6b illustrates this performance enhancement.
@@
C Proof of Impact of Resistor Optimization
To compute the derivatives of $P_{\text{model}}$ with respect to $R_{\text{CoT}}$, we start with the given expression:
$P_{\text{model}}=\frac{\left(\mathcal{E}_{\text{model}}+\mathcal{E}_{\text{ICL}}\right)^{2}R_{0}}{(R_{\text{CoT}}+R_{0})^{2}}.$ (26)
According to Eq. 10, let $R_{\text{CoT}}=R_{1}+R_{2}$, where $R_{1}\gg R_{2}$, if $R_{1}$ is improved to $\frac{R_{1}}{k}$, the power can be calculated as:
$P^{1}_{\text{model}}=\frac{\left(\mathcal{E}_{\text{model}}+\mathcal{E}_{\text{ICL}}\right)^{2}R_{0}}{(\frac{R_{1}}{k}+R_{2}+R_{0})^{2}}$ (27)
@@ C Proof of Impact of Resistor Optimization
$=\frac{\left(\mathcal{E}_{\text{model}}+\mathcal{E}_{\text{ICL}}\right)^{2}R_{0}}{(R_{1}+R_{2}+R_{0})^{2}}-\frac{\left(\mathcal{E}_{\text{model}}+\mathcal{E}_{\text{ICL}}\right)^{2}R_{0}}{(\frac{R_{1}}{k}+R_{2}+R_{0})^{2}}$ (29)
$=\left(\mathcal{E}_{\text{model}}+\mathcal{E}_{\text{ICL}}\right)^{2}R_{0}(\frac{1}{(R_{1}+R_{2}+R_{0})^{2}}-\frac{1}{(\frac{R_{1}}{k}+R_{2}+R_{0})^{2}})$ (30)
@@
$\Delta P^{2}_{\text{model}}=\left(\mathcal{E}_{\text{model}}+\mathcal{E}_{\text{ICL}}\right)^{2}R_{0}(\frac{1}{(R_{1}+R_{2}+R_{0})^{2}}-\frac{1}{(R_{1}+\frac{R_{2}}{k}+R_{0})^{2}})$ (31)
@@
$\Delta P^{1}_{\text{model}}-\Delta P^{2}_{\text{model}}=\left(\mathcal{E}_{\text{model}}+\mathcal{E}_{\text{ICL}}\right)^{2}R_{0}(\frac{1}{(\frac{R_{1}}{k}+R_{2}+R_{0})^{2}}-\frac{1}{(R_{1}+\frac{R_{2}}{k}+R_{0})^{2}})$ (32)
@@
We only need to examine the sign of $B-A$, since $B+A>0$ holds true at all times.since $R_{1}\gg R_{2}$, and $01-\frac{1}{k}>$ 0 for $n>1$, we have:

### v2

@@ B.1.2 Analysis of Effectiveness of Coverage Design
Similarly, we hypothesize that coverage can be considered correct if at least one of multiple parallel reasoning is accurate. Under this assumption, the integration difficulty in self-consistency verification approaches zero. Analogously, this can be modeled as a zero-value combined resistance in parallel with multiple ITR resistors, where the coverage resistance of self-consistency $\hat{R}_{\text{all}}$ is given by:
$\hat{R}_{\text{all}}=R_{0}+R_{S}+\frac{1}{\sum_{i=1}^{n}\frac{1}{R_{\text{ITR}}^{i}}}=R_{0}+\frac{1}{\sum_{i=1}^{n}\frac{1}{R_{\text{ITR}}^{i}}}$ (19)
@@
$P^{\text{cover}}_{\text{out}}=\frac{\left(\mathcal{E}_{\text{model}}+\mathcal{E}_{\text{ITL}}\right)^{2}R_{0}}{(\frac{1}{\sum_{n}\frac{1}{R_{\text{ITR}}}}+R_{0})^{2}},$ (20)
which are much more smaller than $P^{\text{self}}_{\text{out}}$ in the same $n$ sample size with $R_{S}\gg 0$. As the number of samples increases, the curve in Fig. 5c demonstrates that the real performance aligns closely with the theoretical range. Building upon the same bounded assumption of $\{R_{\text{ITR}}^{i}\}_{i=1}^{\infty}$, this setup also defines a resistance lower bound:
$\lim_{n\rightarrow+\infty}\hat{R}_{\text{all}}=R_{0}+\lim_{n\rightarrow+\infty}\frac{1}{\sum_{i=1}^{n}\frac{1}{R_{\text{ITR}}^{i}}}=R_{0}$ (21)
@@
$\hat{P}^{\text{cover}}_{\text{out}}=\frac{\left(\mathcal{E}_{\text{model}}+\mathcal{E}_{\text{ITL}}\right)^{2}}{R_{0}}.$ (22)
As shown in Fig. 5c, the model’s performance gradually converges toward this upper bound and much larger than the performance of self-consistency, demonstrating consistency with the theoretical predictions.
@@
According to the theoretical formulation in Eq. 20, in the reasoning phase where the resistance is significantly greater than $R_{0}$, the overall trend shows that Coverage is approximately proportional to the number of samples. Meanwhile, certain Inference Scaling Laws indicate that Coverage and the logarithm of the number of samples exhibit an approximately proportional relationship. These two observations are not contradictory, as ECP assumes relatively independent sampling. However, we have observed that when the number of samples exceeds 100, an increasing proportion of the samples become duplicates rather than independent samples. Consequently, the probability of generating new samples follows the trend where $n$ samples are expected to yield approximately $\log n$ distinct reasoning samples.
@@ B.2 Analysis of Effectiveness of Fine-grained Self-consistency
$R^{\prime}_{\text{all}}=R_{0}+R_{S}+\frac{1}{\sum_{i=1}^{n}\frac{1}{R_{\text{ITR}}^{i}}}.$ (23)
As shown in Fig. 7a, this process is decomposed into multiple sequential, smaller-scale parallel reasoning and verifications. This refinement transforms a single large parallel reasoning resistance with significant verification resistance into a sequence of smaller parallel reasoning resistance and smaller verification resistance. Formally, this is expressed as:
@@
where $n$ denotes the sampling size, $R_{j}^{S}$ represents the small and fine-grained verification resistance for $j$-th step, and $R_{j}^{i}$ corresponds to smaller reasoning resistance for $j$-th step for $i$-th sampling. The total resistance can be broken down into resistance of sub-step resistor as $R^{i}_{\text{ITR}}=\sum_{j}R_{j}^{i}$.
@@
$R^{\text{fine}}_{\text{all}}-R^{\prime}_{\text{all}}$ $=\sum_{j}R_{j}^{S}-R_{S}+\frac{R_{\text{ITR}}}{n}-\sum_{j}\frac{R_{j}^{i}}{n}$ (25)
@@ B.3 Analysis of Effectiveness of Chain-of-Self-consistency
$R^{\text{CoV}}_{\text{all}}$ $=R_{0}+\frac{1}{\sum_{j=1}^{k}\sum_{i=1}^{n}\frac{1}{R_{\text{S}}}}+\sum_{j=1}^{k}\left(R_{\text{meta}}\right)+\frac{R_{\text{ITR}}}{n}$
$=R_{0}+\frac{R_{\text{S}}}{nk}+kR_{\text{meta}}+\frac{R_{\text{ITR}}}{n}.$
Clearly, as the sample size $n$ and the meta-validation steps $k$ increase, $\frac{R_{\text{S}}}{nk}$ decreases, leading to reduced difficulty in meta-validation. This significantly optimizes the total resistance and improves output power. Fig. 7b illustrates this performance enhancement.
@@
B.4 Proof of Impact of Resistor Optimization
To compute the derivatives of $P_{\text{model}}$ with respect to $R_{\text{ITR}}$, we start with the given expression:
$P_{\text{model}}=\frac{\left(\mathcal{E}_{\text{model}}+\mathcal{E}_{\text{ITL}}\right)^{2}R_{0}}{(R_{\text{ITR}}+R_{0})^{2}}.$ (28)
According to Eq. 11, let $R_{\text{ITR}}=R_{1}+R_{2}$, where $R_{1}\gg R_{2}$, if $R_{1}$ is improved to $\frac{R_{1}}{k}$, the power can be calculated as:
$P^{1}_{\text{model}}=\frac{\left(\mathcal{E}_{\text{model}}+\mathcal{E}_{\text{ITL}}\right)^{2}R_{0}}{(\frac{R_{1}}{k}+R_{2}+R_{0})^{2}}$ (29)
@@ B.4 Proof of Impact of Resistor Optimization
$=\frac{\left(\mathcal{E}_{\text{model}}+\mathcal{E}_{\text{ITL}}\right)^{2}R_{0}}{(R_{1}+R_{2}+R_{0})^{2}}-\frac{\left(\mathcal{E}_{\text{model}}+\mathcal{E}_{\text{ITL}}\right)^{2}R_{0}}{(\frac{R_{1}}{k}+R_{2}+R_{0})^{2}}$ (31)
$=\left(\mathcal{E}_{\text{model}}+\mathcal{E}_{\text{ITL}}\right)^{2}R_{0}(\frac{1}{(R_{1}+R_{2}+R_{0})^{2}}-\frac{1}{(\frac{R_{1}}{k}+R_{2}+R_{0})^{2}})$ (32)
@@
$\Delta P^{2}_{\text{model}}=\left(\mathcal{E}_{\text{model}}+\mathcal{E}_{\text{ITL}}\right)^{2}R_{0}(\frac{1}{(R_{1}+R_{2}+R_{0})^{2}}-\frac{1}{(R_{1}+\frac{R_{2}}{k}+R_{0})^{2}})$ (33)
@@
$\Delta P^{1}_{\text{model}}-\Delta P^{2}_{\text{model}}=\left(\mathcal{E}_{\text{model}}+\mathcal{E}_{\text{ITL}}\right)^{2}R_{0}(\frac{1}{(\frac{R_{1}}{k}+R_{2}+R_{0})^{2}}-\frac{1}{(R_{1}+\frac{R_{2}}{k}+R_{0})^{2}})$ (34)
@@
We only need to examine the sign of BAB-A, since $B+A>0$ holds true at all times.since $R_{1}\gg R_{2}$, and $1-\frac{1}{k}>0$ for $n>1$, we have: